The brake applied to an car reduce the acceleration 6 m/s in the opposite direction to the motion.if the car takes two seconds to stop after the application of the brake.calculate the distance travelled by the car after applying brake
1
Expert's answer
2019-07-12T09:09:26-0400

Let's first find the initial velocity of the car from the kinematic equation:


"v_f = v_i + at,"

here, "v_i" is the initial velocity of the car, "v_f = 0" is the final velocity of the car, "a = -6 m\/s^2" is the deceleration of the car and "t = 2 s" is the time during which the car decelerates.

Then, from this formula we can calculate the initial velocity of the car:


"v_i = -at =- (-6 \dfrac{m}{s^2}) \cdot 2.0 s = 12 \dfrac{m}{s}."

Finally, we can find the distance travelled by the car after applying brake from another kinematic equation:


"d = v_it + \dfrac{1}{2}at^2 = 12 \dfrac{m}{s} \cdot 2.0 s + \dfrac{1}{2} \cdot (-6 \dfrac{m}{s^2}) \cdot (2.0 s)^2 = 12 m."

Answer:

"d =12 m."